# The CMS Phase-2 Fast Beam Condition Monitor prototype test with beam

### v1

@@ 1.1 Motivation and requirements for the dedicated CMS luminosity detector
The Fast Beam Condition Monitor (FBCM) is a standalone luminometer designed for precision luminosity measurements and beam-induced background (BIB) [1] monitoring that exploits silicon (Si) pad sensors with a fast, untriggered readout. It is part of the upgrade of the CMS experiment for the High-Luminosity LHC (HL-LHC) era, developed within the Beam Radiation, Instrumentation, and Luminosity (BRIL) project. A precise luminosity measurement is critical for numerous CMS measurements, and the target precision is set to about 2% in real time (online) and below 1% after data post-processing (offline) [1].
Response linearity and stability are two key aspects which define the design of a luminometer. The linearity of FBCM was studied and optimized in simulations [1] for an average rate of interactions per bunch crossing (also referred to as pileup) up to . Linearity across a wide pileup range is essential as the specialized calibration runs are conducted at a low pileup of a few times per year, while standard physics data taking in proton-proton collisions during HL-LHC operations will occur at . The stability of the luminometer is influenced by operational aspects, as well as the aging of the sensors and electronics under radiation exposure.
The simple and modular design of FBCM uses radiation-hard components for efficient operation until its scheduled replacement in the middle of the HL-LHC period. The expected total ionizing dose (TID) at that time is estimated to be approximately 80 MRad at the location of the Si-pad sensors at a radius of cm, corresponding to a 1 MeV neutron equivalent fluence of about and an integrated luminosity of about 2000 .
@@ 1.2 FBCM detector overview
A first design of the FBCM was presented in 2021 in the BRIL Phase-2 Technical Design Report [1]. Since then, extensive development work was performed to improve, validate, and finalize the design. It lead to the (pre-)production and testing of a dedicated front-end ASIC [2] and various printed circuit boards (PCB) for prototyping the individual components [3], the finalization of the readout, control and powering systems, as well as the cabling plan, and the production of a real size mock-up [4].
@@ 1.2.2 FBCM front-end electronics
A list of all the test boards which are described in the paper with tracking number and important characteristics is presented in Table 1. For the boards with irradiated sensors and ASICs, the approximate total ionizing dose is given. The TID ranges presented for the irradiated sensors are broad, as the proton beam used for irradiation did not cover the sensors uniformly, causing variations across the pads of a single sensor. In contrast, the ASICs were irradiated under a wide X-ray beam [8], resulting in a more precise TID.
@@ 1.2.3 FBCM silicon sensors
There are two types of sensors available for FBCM, as shown in Fig. 4. The main difference between them is the thickness and the guard ring structure, which is required to restrict the active area of the sensor. The active area per pad, defined by the size of the metallization, is the same for both sensor options — each pad measures 1.7 x 1.7 mm2. The FBCM ASIC is designed to be compatible with both types of sensor. However, shot noise increases with increasing leakage current under accumulated sensor exposure to radiation. The initial ASIC noise is 650 e- [2], but after exposure to cm-2 1 MeV neutron equivalent () fluence, it increases to about 1500 e- for the 290 m thick two-pad sensors and 850 e- for 150 m thick six-pad sensors. The expected signal in electrons for a particular sensor can be estimated by multiplying the sensor thickness by the number of electron-hole pairs produced in silicon =78 m-1 [9]. The expected signal-to-noise (S/N) ratio and other characteristics of the sensors, both before and after irradiation, are listed in Table 2, taking into account the reduced charge collection efficiency (CCE) of the sensors after irradiation [10]. The capacitance is given based on the simple reverse-biased diode capacitance model.
@@
Leakage current after irradiation at -20 at 500 V | 1 | 0.1 |
@@
Two-pad sensors were previously used in CMS for the BCM1F detector [11] during Run 3 of the LHC and exhibited current instability after exposure to a fluence of about cm-2 1 MeV neutron equivalent. This instability [12] was attributed to the build up of surface current and the formation of a conductive path, which allowed current to flow from the edges of the sensor to the guard ring. To reproduce this behavior under controlled laboratory conditions, the two-pad sensors were tested under an X-ray lamp [8], which generates only surface charge. Over ten hours of X-ray exposure, a slow increase of the current was observed. When the X-ray lamp was switched off, a gradual current decrease was observed, lasting about one hour. To verify that the double guard ring structure of the six-pad sensor resolves this issue, the same test was repeated. The results showed that the current through the guard ring and DC-pads remained stable after hours of the surface irradiation, making the six-pad sensor the preferred choice for FBCM construction. However, both types of sensors were thoroughly tested in the proton beam test in July 2024.
@@ 3.1 Time over threshold
Fig. 17 (right) shows the effect of RC3 versus RC6 ASIC settings on the ToT histogram for the same input channel and at a fixed 0.9 fC threshold value. As the shaping time of the signal is different for these two settings, the expected shift of the ToT to higher value for RC6 is observed. This is beneficial for irradiated sensors, as it compensates for the lower charge collection efficiency, but has to be operated in combination with higher thresholds to cut the noise. The integrals of the histograms aggregated for the same number of spills with RC3 and RC6 settings agree within 5%.
@@ 3.2 Time walk
The ToA is affected by the input charge due to “time walk": the lower the input charge, the later the threshold is reached. Time walk was measured in the laboratory and defined as a delay between the trigger initiating the injection of a fixed-charge calibration pulse and the detected arrival time of the pulse, shifted such that the time walk is zero at maximum charge injection. Calibration pulses from 1 to 12.5 fC charge were tested. For an injected charge of 1.9 fC, as expected in a sensor with a width of 150 m (assuming 78 electron-hole pairs per m created by a MIP crossing the silicon sensor [9]), the measured time walk is about 2 ns, as shown with blue dots in Fig. 18 (left). This matches the FBCM23 ASIC specifications [2]. At an input charge of 1 fC, the delay reaches about 5 ns. On this scale the time walk does not affect the beam induced background measurements or per bunch crossing luminosity measurements, as bunch spacing is 25 ns.
@@ 3.3 Charge deposition calibration
The difference between the expected 150 thickness of a six-pad sensor and the calculated effective thickness lies within 9% (using the logarithmic fit) and 2% (using the linear interpolation) for this example. However, for other channels, boards and RC filter settings, the difference could be as large as 30%, as indicated in Table 3. The discrepancy is caused by several factors, such as the uncertainty in the approximation of the calibration curve or the difference between the signal shape of the injected charges during calibration and the actual shape induced by charges moving through the Si-sensor, with the latter having a larger time constant, especially for irradiated sensors.
@@ 3.4 Effects of irradiation on time over threshold
For some channels corresponding to the irradiated six-pad sensors (board 21), there is a significant difference between the expected shape described by a convolution of a Landau and a Gaussian function and the measured shape of the ToT distribution. This is most evident for channels 0 and 3 of board 21, where a “shoulder” is noticeable at lower ToT values. More pronounced “tails” than expected are also visible for all channels in all boards at higher ToT values. This deviation from the ideal shape varies depending on the threshold and is studied in more detail using the track pointing information described in Sec. 3.8 and with the transient current technique described in Sec. 3.9.
@@ 3.7 Triggered and untriggered readout
FBCM is designed for untriggered operation, however, as a cross–check during the beam test, a triggered readout was implemented. Details of the triggered readout are given in Sec. 2.2. At a low threshold of 0.3 fC, where significant contribution from noise (ToT values below 5 ns) is expected, the triggered ToT distribution differs from the untriggered one due to the efficient noise suppression of the hit requirement in the active region of the SCC, as shown in Fig. 25 on the left. With the threshold increased to 0.5 fC, as shown in Fig. 25 on the right, the noise is suppressed and the two distributions are almost indistinguishable.
The presence of two signal peaks in the measured ToT distribution is unexpected. It was reproduced with very low intensity electron and muon beams to verify that the effect is not dependent on the particle type or beam intensity. It was also investigated later in the laboratory with controlled laser pulse injections (see Sec. 3.9).
@@
Fig. 26 (left) shows the ToT distribution for an irradiated two-pad sensor. A small peak with a mean of around 9 ns is noticeable. This peak was much less pronounced for unirradiated two-pad sensors and is not visible for ASIC thresholds above 0.9 fC (see red line in Fig. 20). To study this effect, the ToT distribution for tracks associated with the inner region of the sensor were plotted separately from those associated with the border region. The definitions of the border and inner regions of the sensor are discussed in Sec. 3.6. From Fig. 27 (left), it is clear that tracks associated with the border region of the sensor (red) have a dominant contribution to the peak at low ToT values (below 13 ns).
@@ 3.9 Transient current technique study of six-pad sensors
An example of the ToT histogram collected for a low ASIC threshold of 0.3 fC is shown in Fig. 30 on the left and individual waveforms are shown on the right. Short noise pulses accumulate in the first peak of the ToT histogram, below 5 ns. The second peak is associated with laser injection, but it has a substructure. It results from the degraded timing response caused by the slow traveling charges, which affect the ToT and temporarily lower the discriminator’s effective threshold following the prompt signal generated by the collected electrons. This temporary threshold reduction increases the front-end’s susceptibility to after-pulsing, which may be induced by either preamplifier electronic noise or electromagnetic interference (EMI), i.e. electronic pickup. Electronic pickup noise is setup-dependent and not related to the ASIC or the sensor. The time window affected by the after-pulsing is captured between the blue dotted lines in Fig. 29 (right). Noise pulses with arrival time in this time window (red dotted line) cause longer ToT, shown in region 2 of Fig. 30 (left). The same mechanism and coincidence with a longer noise pulse correspond to region 3 of Fig. 30 (left). If no noise pulse arrives within the window, the "nominal" ToT is observed in region 1.
@@ 4 Prospects and future work
The beam test results provide new prospects for tuning of the FBCM ASIC simulation based on real data for different ASIC parameters. In simulation, the input pulse shape can be varied to achieve the best agreement with the beam test results for fresh and irradiated sensors of different thicknesses. Then, this tuned simulation can be used to fill in the ASIC parameter space between the beam test measurements. For example, the simulation of the ToT spectra for intermediate irradiation and all possible RC settings should allow for an optimal choice of RC setting for different eras of detector operation, depending on the aggregated dose.
Simulation can be also used to study possible oversampling options for the ToA distribution. In particular, to show if there is a significant ToA improvement in using 8 or 10 samples per bunch crossing, compared to 4 and 6 samples used in Run 3 for the BCM1F detector. Results of this study will be used to motivate the choice of the ToA sampling in the FBCM firmware.
@@ 5 Summary
The design of the test board (which also served as the very first module prototype) has been optimized by replacing the pitch adapter with direct bonding between the sensor pads and the FBCM23 ASIC inputs to reduce pickup noise. The effect of the pickup noise on the time over threshold distribution was studied with the transient current technique, and an optimal bias voltage and ASIC threshold for FBCM operation were proposed. The current design of the test board with a 150 m thick six-pad sensor with double guard rings and two FBCM23 ASICs was approved after the beam test and is being used as the starting point for the front-end module design which is expected to be ready and prototyped by summer 2025, in time for the next beam test with (close to) final front- and back-end components.

### v2

@@ 1.1 Motivation and requirements for the dedicated CMS luminosity detector
The Fast Beam Condition Monitor (FBCM) is a standalone luminometer designed for precision luminosity measurements and beam-induced background (BIB) [1] monitoring that exploits silicon (Si) pad sensors with a fast, untriggered readout. It is part of the upgrade of the CMS experiment for the High-Luminosity LHC (HL-LHC) era, developed within the Beam Radiation, Instrumentation, and Luminosity (BRIL) project. A precise luminosity measurement is critical for numerous CMS measurements, and the target precision on the luminosity is set to about 2% in real time (online) and below 1% after data post-processing (offline) [1].
During HL-LHC wide range of an average rate of interactions per bunch crossing (also referred to as pileup) must be covered: from up to . Linearity of the luminometer response for the full range and its stability are two key aspects which define the design of a luminometer. Linearity across a wide pileup range is essential as the specialized calibration runs are conducted at a low pileup of a few times per year, while standard physics data taking in proton-proton collisions during HL-LHC operations will occur at . The linearity of FBCM was studied and optimized in simulations [1]. Detector is shown to be linear up to withing uncertainty band. Stability of the detector is defined as ability to maintain its metrological properties, such as accuracy and performance, over extended period of time (on the scale of several years for luminometer). The stability of the luminometer is influenced by operational aspects, as well as the aging of the sensors and electronics under radiation exposure.
The simple and modular design of FBCM uses radiation-hard components for efficient operation until its scheduled replacement in the middle of the HL-LHC period. The expected total ionizing dose (TID) at that time is estimated to be approximately 80 MRad at the location of the Si-pad sensors at a radius of cm, corresponding to a 1 MeV neutron equivalent fluence of about and an integrated luminosity of about 2000 (see Fig 3.1 and Fig.18.12 in [1]). The mixed radiation field at this location is composed of various particle types with a charged lepton to charged hadron to neutron fluence ratio of approximately 0.5: 0.75: 1.
@@ 1.2 FBCM detector overview
A first design of the FBCM was presented in 2021 in the BRIL Phase-2 Technical Design Report [1]. Since then, extensive development work was performed to improve, validate, and finalize the design. It led to the (pre-)production and testing of a dedicated front-end ASIC [2] and various printed circuit boards (PCB) for prototyping the individual components [3], the finalization of the readout, control and powering systems, as well as the cabling plan, and the production of a real size mock-up [4].
@@ 1.2.2 FBCM front-end electronics
A list of all the test boards which are described in the paper with tracking number and important characteristics is presented in Table 1. For the boards with irradiated sensors and ASICs, the approximate total ionizing dose is given. The TID ranges presented for the irradiated sensors are broad, as the 24 GeV proton beam used at CERN IRRAD facilities [8] did not cover the sensors uniformly, causing variations across the pads of a single sensor. In contrast, the ASICs were irradiated under a wide X-ray beam [9], resulting in a more precise TID.
@@ 1.2.3 FBCM silicon sensors
There are two types of n-on-p sensors available for FBCM, as shown in Fig. 4. Both sensor types use float zone thinned (FTH150) wafer material [5, 10] and all sensors are produced by a single vendor, Hamamatsu Photonics K.K. (HPK). The main difference between them is the thickness and the guard ring structure, which is required to restrict the active area of the sensor. The active area per pad, defined by the size of the metallization, is the same for both sensor options — each pad measures 1.7 x 1.7 mm2. The FBCM ASIC is designed to be compatible with both types of sensor. However, shot noise increases with increasing leakage current under accumulated sensor exposure to radiation. The initial ASIC noise is 650 e- [2], but after exposure to cm-2 1 MeV neutron equivalent () fluence, it increases to about 1500 e- for the 290 m thick two-pad sensors and 850 e- for 150 m thick six-pad sensors. The expected signal in electrons for a particular sensor can be estimated by multiplying the sensor thickness by the number of electron-hole pairs produced in silicon =78 m-1 [11]. The expected signal-to-noise (S/N) ratio and other characteristics of the sensors, both before and after irradiation, are listed in Table 2, taking into account the reduced charge collection efficiency (CCE) of the sensors after irradiation [12]. The capacitance is given based on the simple reverse-biased diode capacitance model.
@@
| Measured leakage current after irradiation at -20 at 500 V | 1 | 0.1 |
@@
Two-pad sensors were previously used in CMS for the BCM1F detector [13] during Run 3 of the LHC and initially operated at bias voltage of 400 V. As they exhibited current instability after exposure to a fluence of about cm-2 1 MeV neutron equivalent, operational bias voltage was lowered to 350 V. This instability [14] manifested itself in the unpredicted rise of the leakage current and eventual trip due to reached set power supply current limit (50 A); it was attributed to the build up of surface current and the formation of a conductive path, which allowed current to flow from the edges of the sensor to the guard ring. Reduction of the bias voltage reduces the frequency of the trips, but did not mitigate it. Further reduction was not possible due to loose off sensors efficiency. To reproduce this behavior under controlled laboratory conditions, the two-pad sensors were tested under an X-ray lamp [9], which generates only surface charge. Over ten hours of X-ray exposure, a slow increase of the current was observed. Intensity of X-ray was selected to match the leakage current close to that observed in operations at the beginning of collisions (10 A). When the X-ray lamp was switched off, a gradual current decrease was observed, lasting about one hour. To verify that the double guard ring structure of the six-pad sensor resolves this issue, the same test was repeated. The results showed that the current through the guard ring and DC-pads remained stable after hours of the surface irradiation, making the six-pad sensor the preferred choice for FBCM construction. However, both types of sensors were thoroughly tested in the proton beam test in July 2024.
@@ 3.1 Time over threshold
Fig. 17 (right) shows the effect of RC3 versus RC6 ASIC settings on the ToT histogram for the same input channel and at a fixed 0.9 fC threshold value. As the shaping time of the signal is different for these two settings, the expected shift of the ToT to higher value for RC6 is observed. This is beneficial for irradiated sensors, as it compensates for the lower charge collection efficiency, but has to be operated in combination with higher thresholds to cut the noise. In this particular example utilizing not irradiated sensor, RC6 integral is about 15% lower than that of RC3, indicating efficiency drop. While it was checked that for the irradiated sensors the integrals of the histograms aggregated for the same number of spills with RC3 and RC6 settings agree within 5%.
@@ 3.2 Time walk
The ToA is affected by the input charge due to “time walk": the lower the input charge, the later the threshold is reached. Time walk was measured in the laboratory and defined as a delay between the trigger initiating the injection of a fixed-charge calibration pulse and the detected arrival time of the pulse, shifted such that the time walk is zero at maximum charge injection. Calibration pulses from 1 to 12.5 fC charge were tested. For an injected charge of 1.9 fC, as expected in a sensor with a width of 150 m, the measured time walk is about 2 ns, as shown with blue dots in Fig. 18 (left). This matches the FBCM23 ASIC specifications [2]. At an input charge of 1 fC, the delay reaches about 5 ns. On this scale the time walk does not affect the beam induced background measurements or per bunch crossing luminosity measurements, as bunch spacing is 25 ns.
@@ 3.3 Charge deposition calibration
The difference between the expected 150 thickness of a six-pad sensor and the calculated effective thickness lies within 9% (using the logarithmic fit) and 2% (using the linear interpolation) for this example. However, for other channels, boards and RC filter settings, the difference could be as large as 30%, as indicated in Table 3. Higher deviation is observed for the board 21 with irradiated sensors. The discrepancy is caused by several factors: the uncertainty in the approximation of the calibration curve; the difference between the signal shape of the fast injected charges directly to the ASIC input during calibration and the actual shape induced by charges moving through the Si-sensor. The latter having a larger time constant, especially for irradiated sensors explains why observed difference from the expected thickness increases for irradiated sensors.
@@ 3.4 Effects of irradiation on time over threshold
For some channels corresponding to the irradiated six-pad sensors (board 21), there is a significant difference between the expected shape described by a convolution of a Landau and a Gaussian function and the measured shape of the ToT distribution. This is most evident for channels 0 to 3 of board 21, where a “shoulder” is noticeable at lower ToT values. More pronounced “tails” than expected are also visible for all channels in all boards at higher ToT values. This deviation from the ideal shape varies depending on the threshold and is studied in more detail using the track pointing information described in Sec. 3.8 and with the transient current technique described in Sec. 3.9.
@@ 3.7 Triggered and untriggered readout
FBCM is designed for untriggered operation, however, as a cross–check during the beam test, a triggered readout was implemented. Details of the triggered readout are given in Sec. 2.2. At a low threshold of 0.3 fC, where significant contribution from noise (ToT values below 5 ns) is expected, the triggered ToT distribution differs from the untriggered one, as shown in Fig. 25 on the left. This efficient noise suppression is achieved requiring signal from the auxiliary detector (active mask of the SCC, see Sec. 2.2). With the threshold increased to 0.5 fC, the noise is already suppressed by the threshold, no additional suppression is added by triggered readout, and the two distributions are almost indistinguishable, as shown in Fig. 25 on the right.
The presence of two signal peaks in the measured ToT distribution is unexpected. From the expected energy deposition only the peak around ToT of 10 ns was expected. However, double peak structure was reproduced with very low intensity electron and muon beams to verify that the effect is not dependent on the particle type or beam intensity. It was also investigated later in the laboratory with controlled laser pulse injections (see Sec. 3.9).
@@
Fig. 26 (left) shows the ToT distribution for an irradiated two-pad sensor. A small peak with a mean of around 9 ns is noticeable. This peak was much less pronounced for unirradiated two-pad sensors, presumably due to less defects and more homogeneous electric field at the edges, and is not visible for ASIC thresholds above 0.9 fC (see red line in Fig. 20). To study this effect, the ToT distribution for tracks associated with the inner region of the sensor were plotted separately from those associated with the border region. The definitions of the border and inner regions of the sensor are discussed in Sec. 3.6. From Fig. 27 (left), it is clear that tracks associated with the border region of the sensor (red) have a dominant contribution to the peak at low ToT values (below 13 ns).
@@ 3.9 Transient current technique study of six-pad sensors
An example of the ToT histogram collected for a low ASIC threshold of 0.3 fC is shown in Fig. 30 on the left and individual waveforms are shown on the right. Short noise pulses accumulate in the first peak of the ToT histogram, below 5 ns. The second peak is associated with laser injection, but it has a substructure. It results from the degraded timing response caused by the slow traveling charges, which affect the ToT and temporarily lower the discriminator’s effective threshold following the prompt signal. This temporary threshold reduction increases the front-end’s susceptibility to after-pulsing, which may be induced by either preamplifier electronic noise or electromagnetic interference (EMI), i.e. electronic pickup. Electronic pickup noise is setup-dependent and not related to the ASIC or the sensor. The time window affected by the after-pulsing is captured between the blue dotted lines in Fig. 29 (right). Noise pulses with arrival time in this time window (red dotted line) cause longer ToT, shown in region 2 of Fig. 30 (left). The same mechanism and coincidence with a longer noise pulse correspond to region 3 of Fig. 30 (left). If no noise pulse arrives within the window, the "nominal" ToT is observed in region 1.
@@ 4 Prospects and future work
The beam test results provide new prospects for tuning of the FBCM ASIC simulation based on real data for different ASIC parameters. In simulation, the input pulse shape can be varied to achieve the best agreement with the beam test results for fresh and irradiated sensors of different thicknesses. Then, this tuned simulation can be used to populate in the ASIC parameter space where direct beam test measurements are missing. For example, the simulation of the ToT spectra for intermediate irradiation and all possible RC settings should allow for an optimal choice of RC setting for different eras of detector operation, depending on the aggregated dose.
As final FBCM deliverable used for luminosity is ToA measurement, more studies of the ToA have to be done. In particular, a decision must be taken on the number of samples per colliding bunch. Simulation can be also used to study possible oversampling options for the ToA distribution. In particular, to show if there is a significant ToA improvement in using 8 or 10 samples per bunch crossing, compared to 4 and 6 samples used in Run 3 for the BCM1F detector. Results of this study will be used to motivate the choice of the ToA sampling in the FBCM firmware.
@@ 5 Summary
The design of the test board (which also served as the very first module prototype) has been optimized by replacing the pitch adapter with direct bonding between the sensor pads and the FBCM23 ASIC inputs to reduce pickup noise. The effect of the pickup noise on the time over threshold distribution was studied with the transient current technique, and an optimal bias voltage and ASIC threshold for FBCM operation were proposed. The current design of the test board with a 150 m thick six-pad sensor with double guard rings and two FBCM23 ASICs was approved after the beam test and is being used as the starting point for the front-end module design. Stability of the new six-pad sensor was tested with X-ray, and despite shorter efficiency vs. threshold range (due to smaller sensors thickness) was prioritized above two-pad sensor. New front-end is expected to be ready and prototyped by summer 2025, in time for the next beam test with (close to) final front- and back-end components.